OpenStudy (anonymous):

Find the sum to n terms of the series : \[\frac{1}{2}+\frac{3}{4}+\frac{7}{8}+\frac{15}{16}+.........\]

mathslover (mathslover):

oh k so first of all it is an AP or GP or HP ... ? @shubham.bagrecha can u tell me?

OpenStudy (anonymous):

i can be converted into A.G.P by taking 1/2 as x.

mathslover (mathslover):

good .. so by taking 1/2 as x we have: \[\large{x+ \frac{3x}{2} + \frac{7x}{4}+...}\]

mathslover (mathslover):

right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[x + 3x^{2}+7x^{3}+15x^{4}+...............\]

mathslover (mathslover):

hmn we can write like that also wait ..

OpenStudy (anonymous):

hint :\[a_n=\frac{2^n-1}{2^n}\]

OpenStudy (anonymous):

this is for GP only ??

OpenStudy (anonymous):

@mukushla ??

OpenStudy (anonymous):

no this is the whole thing..

OpenStudy (anonymous):

u can break it into\[\frac{2^n-1}{2^n}=1-(\frac{1}{2})^n\]

OpenStudy (anonymous):

second part is GP

OpenStudy (anonymous):

u can sum up it into n tems now

OpenStudy (anonymous):

*terms

OpenStudy (anonymous):

how this formula came?

OpenStudy (anonymous):

can you explain step-wise.

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{3}{4}+\frac{7}{8}+\frac{15}{16}+...=\frac{2-1}{2}+\frac{4-1}{4}+\frac{8-1}{8}+\frac{16-1}{16}+...\\=\frac{2^1-1}{2^1}+\frac{2^2-1}{2^2}+\frac{2^3-1}{2^3}+\frac{2^4-1}{2^4}+...\]

OpenStudy (anonymous):

we are told to substitute 1/2 for x , then how will we solve?

OpenStudy (anonymous):

for GP part right?

OpenStudy (anonymous):

x+3x^2+7x^3+15x^4+...............

OpenStudy (anonymous):

is this a GP ? this is not

OpenStudy (anonymous):

how can we sum it into n term?

OpenStudy (anonymous):

there ia a GP along with AP whose difference is also in GP

OpenStudy (anonymous):

@hartnn

mathslover (mathslover):

@TuringTest

hartnn (hartnn):

i m on 2 other questions right now,after that i will look through this...

hartnn (hartnn):

let me give u the formula for arithmetico-geometric sequence defined as:\[x_n=a_ng_n\] , where an and gn are the 'n'th terms of arithmetic and geometric sequences, respectively. \[\frac{a_{n}g_{n+1}}{r-1}-\frac{x_{1}}{r-1}-\frac{d(g_{n+1}-g_{2})}{(r-1)^{2}} \] or \[ \frac{a_{n}g_{n+1}-x_{1}-drS_{g}}{r-1} \] d is common difference of an and r is common ratio of gn,x1 if 1st term of your AGP or equivalently u can use \[t_{n}=\left[a+\left(n-1\right)d\right]r^{n-1}\] and sum Sn will be: \[S_{n}=\frac{a}{1-r}+\frac{dr\left(1-r^{n-1}\right)}{(1-r)^{2}}-\frac{\left[a+\left(n-1\right)d\right]r^{n}}{1-r}\]